There are 8 yellow and 6 blue marbles. How many yellow marbles can you add to make the ratio of yellow to blue 2:1
Answer by swincher4391(1107)

So currently we have 8 yellow and 6 blue marbles. We need to add an amount to yellow marbles to make it so that the ratio is 2:1.
We can set it up like this. Let x be the amount that we add to the yellow marbles.
Then (8+x)/6 = 2/1
So 8+x = 12 by cross multiplication.
x = 4
So we can add 4 yellow marbles to get a ratio of 2:1.